OpenStudy (anonymous):

solve using the sustitution method x=y-1 y=4-2x

OpenStudy (anonymous):

y=4 -2(y-1) y=-2y+6 y=2 x=1

OpenStudy (anonymous):

\[y=4-2(y-1)\]

OpenStudy (anonymous):

ankur beat me to it..but what he has is right.

OpenStudy (anonymous):

thank you